OpenStudy (anonymous):

I need help with this problem. Solve the proportion 2x/5x-6 = 2/x.. What is the value of x?

OpenStudy (anonymous):

\[\frac{ 2x }{ 5x -6 }=\frac{ 2 }{ x }\]